4. What is the pressure (psia) at the base of a 8 ft liquid column with atmospheric pressure 755 mm Hg, if

the fluid is a) water, b) glycerin (SG = 1.25), c) Oil (SG = 0.750)


1
Expert's answer
2021-09-08T07:44:17-0400

"P =P_{\text{atm}}+\rho gh"


"P_{\text{atm}} =\dfrac{755\text{ mmHg}}{760\text{ mmHg}}× 101325\ Pa"

"=100658 Pa"

"h = 8\ ft= 2.44\ m"


a) "P =P_{\text{atm}}+\rho gh\\
P = 100658 + 1000(9.8)(2.44) = 124570\ Pa""P =P_{\text{atm}}+\rho gh\\
P = 100658 + 1000(9.8)(2.44)(1.25) =\\"


b)"P =P_{\text{atm}}+\rho gh\\
P = 100658 + 1000(9.8)(2.44)(1.25)\\= 130548
\ Pa"


c)

"P =P_{\text{atm}}+\rho gh\\
P = 100658 + 1000(9.8)(2.44)(0.75)\\ = 118592
\ Pa"